A painter is painting a sign on a bridge. A can of paint falls 10m to the ground below. Use the Conservation of Energy to calculate the velocity with which the can hits the ground. Give your answer correct to 2 decimal places. Clearly state your unit.

Take g = 10m/s2
Answer #1

Applying conservation of energy,

Potential energy= kinetic energy

mgh = 1/2 ×mv2

v = √2gh

v =√2×10×10 = √200

v =14.14 m/s